# Generative AI Enables EEG Super-Resolution via Spatio-Temporal Adaptive Diffusion Learning

### v4

@@ I Introduction
Electroencephalogram (EEG) is a economical, non-invasive neuroimaging technology with high temporal resolution, playing a vital role in cognitive neuroscience research and clinical diagnosis. Unlike other neuroimaging techniques such as positron emission tomography (PET) and functional magnetic resonance imaging (fMRI), EEG offers superior temporal resolution and can continuously monitor brain activity in natural environments. This capability substantially supports the study of dynamic cognitive processes [1]. Recently, the development of portable and wireless EEG devices has further improved the flexibility and usability of EEG, enabling its widespread application in emerging neural engineering scenarios such as epileptic seizure detection and brain-computer interface (BCI) control [2, 3, 4]. Besides, these consumer devices can realize real-time remote monitoring of patients through the Internet of Medical Things (IoMT), enhance the early diagnosis and treatment, and improve the performance of real-time applications in intelligent consumer healthcare systems [5]. However, the devices typically have fewer than 64 channels, resulting in lower spatial resolution of the acquired EEG data, which severely limits its potential in clinical diagnosis and brain function imaging research [6]. Currently, high-density (HD) EEG devices improve the spatial resolution of EEG by employing more electrode quantities, thus assisting doctors in performing evaluations or diagnoses in different clinical scenarios. For example, recent studies have shown that non-invasive HD EEG devices can objectively, non-invasively, and accurately determine the epileptic focus area [7, 8, 9]. However, HD EEG devices greatly increase hardware and acquisition costs, their bulkiness and complex wiring prevent subjects from moving their heads freely, restricting their use to controlled environments such as laboratories or hospitals. In addition, due to poor wearing comfort, subjects cannot be monitored for extended periods [10]. Fig. 1 summarizes the respective advantages and limitations of different density EEG devices and reveals difficulties in acquisition and limited application scenarios for HD EEG devices.
@@
Currently, in the field of enhancing spatial resolution of EEG, the methodologies predominantly fall into two main categories: channel interpolation based on mathematical models and EEG SR reconstruction utilizing deep learning methods. Channel interpolation methods principally rely on predefined scalp surface models to reconstruct missing channel data using signals from other channels, thereby improving spatial resolution. These methods involve calculating the second-order derivatives of the interpolation function in space to explore the distribution across the electrode channels [13]. Subsequent studies have attempted more precise measurements of the spatial distances between EEG electrodes and more accurate predictions for missing channel data, for example, using inverse distance weighting based on Euclidean distances and multi-quadratic interpolation methods. However, these methods only utilize the spatial correlations among adjacent electrodes and typically struggle to effectively capture and reconstruct the high-frequency and transient components of brain electrical activity, resulting in suboptimal reconstruction performance for non-stationary EEG signals.
@@
Lately, deep learning has become the mainstream method in multiple fields, including the field of Internet healthcare [14, 15, 16] and medical image processing [17, 18]. Specifically, deep learning-based EEG SR technology involves training deep learning models to learn the implicit mappings between LR EEG and HR EEG, and using these trained models to reconstruct SR EEG. Notably, research has introduced EEG SR techniques based on convolution neural networks (CNN) [19], followed by studies proposing deep EEG SR frameworks using graph convolution networks (GCNs) [20] to correlate the structural and functional connections of participants’ brains [21]. Additionally, generative models offer a new learning paradigm for medical image reconstruction [22]. Existing researches have utilized approaches such as Autoencoders (AE) and Generative Adversarial Networks (GANs) [23] to enhance the spatial resolution of EEG. Although current EEG SR methods based on deep learning have made considerable advancements in enhancing EEG spatial resolution, they are still limited in effectively capturing both the temporal dynamics and spatial features of EEG, resulting in limited generalizability and practicality of the models. Moreover, the enhanced spatial resolution achieved by these methods is constrained, with a maximum capability of reconstructing EEG data up to 64-channel level, which may not meet the requirements of specific clinical scenarios.
@@
In clinical diagnostic and therapeutic settings, achieving high-fidelity reconstruction of HR (256 channels) ground truth EEG presents a substantial challenge. The key is accurately depicting the dependencies between LR EEG and HR EEG and addressing the significant channel-level disparity between them. Recently, a novel generative model—diffusion models [24] have exhibited remarkable generative capabilities in computer vision, leading to a significant increase in the Artificial Intelligence Generated Content (AIGC) [25, 26]. Moreover, diffusion models have demonstrated superior image generation quality compared to GANs, while also providing advantageous features such as comprehensive distribution coverage, extensibility, and stable training targets [27]. Besides, Latent Diffusion Models (LDMs) [28] extend diffusion models into latent spaces, utilizing cross-attention mechanisms to achieve refined control in conditional generation tasks. Furthermore, SynDiff [29] and DiffMDD [30] have illustrated that diffusion models are applicable in generating medical images, thereby advancing the accuracy of disease diagnosis. Despite this, Applying diffusion models to enhance EEG spatial resolution has not been thoroughly explored. Therefore, this paper aims to investigate the potential of diffusion models in achieving EEG SR, particularly in enhancing spatial resolution. Through this novel generative model, we hope to provide a novel and effective method for processing and analyzing EEG data, supporting more accurate EEG studies and applications across various neuroscience contexts.
@@ III-B Basic ideas
As Fig. 2 shows, the entire model framework consists of EEG pre-trained autoencoder, STC and MTD. During the training phase of STAD, the HR EEG ground truth is first input into the EEG pre-trained encoder to obtain the corresponding latent vectors $z_{0}$. The forward diffusion process progressively derives the state $z_{t}$ of each time step $t$. In the reverse denoising process, the STC first encodes the time series and channel spatial distribution matrix from LR EEG into the spatio-temporal representation conforming to the subject, and concatenates it with time embedding as the conditioning vector. MTD takes the conditioning vectors and noise sampled from the standard normal distribution as input, implements the conditional guidance mechanism through the cross-attention layer, and predicts the noise at each step of the reverse process. To train the MTD, the Mean Squared Error (MSE) loss is minimized between the predicted and actual noise at each step, ultimately generating SR EEG through progressive denoising. During the sampling phase of STAD, the deterministic MTD precisely generates SR EEG matching the input LR EEG by progressively sampling using the spatio-temporal conditional information of LR EEG and noise sampled from the Gaussian distribution.
@@ III-C2 Spatio-temporal Condition Module
As the yellow part of Fig. 2 shows, the STC is mainly composed of a spatial position embedding layer, a 1D convolution block, and a Transformer block. More specifically, the channel spatial embedding layer encodes the channel spatial coordinates of LR EEG, outputting a set of spatial features. Then, we partition the input time series into different patches and process them with the 1D convolution block for the initial extraction of temporal features [43]. Finally, the Transformer block concatenates the temporal features with position encoding and channel spatial embeddings to output the final spatio-temporal condition $c$. The 1D convolution block consists of a 1D convolution layer, batch normalization layer and a ReLU activation function. In the Transformer block, a multi-head self-attention (MSA) layer is followed by feed-forward networks, both of which are encapsulated with residual connections and layer normalization.
@@ III-C3 Multi-scale Transformer Denoising Module
EEG, limited by the number and placement of electrodes, struggles with spatial resolution. High-density EEG devices have shown promise in tasks like source localization and BCI. The multi-scale Transformer denoising module is designed to enhance the spatial resolution of existing EEG devices. Traditional diffusion models, which typically use UNet, are effective in image generation but limited in handling the long-sequence dependencies of time-series data like EEG. Inspired by [39], the MTD employs a Transformer backbone to more effectively extract spatio-temporal features. Additionally, multi-scale 1D convolution blocks are introduced to capture multi-scale temporal features in the reverse denoising process.
@@ III-D Loss Function
where $\epsilon_{\theta}$ is the proposed multi-scale Transformer denoising model and $\tau_{\theta}$ represents the proposed STC. Algorithm 1 presents the entire training procedure along with the final loss function.
@@
      Perform: Extract spatio-temporal features of LR EEG $x$ via STC ;
      $c=STC\left(x\right)$ ;
@@ IV-A Dataset and Preprocessing
We employed the MNE software for preprocessing of the raw data. In this work, we applied high-pass filtering to the continuous EEG time series to remove low-frequency irrelevant information. To eliminate the large line noise caused by direct current, we used notch filtering at 50, 100, 150, and 200 Hz. Subsequently, we segmented the raw data into epochs of fixed duration (350 ms) based on the provided annotation events in the dataset. Moreover, the dataset was recorded using the EGI NA-400 amplifier (Electrical Geodesics) [48] at a sampling rate of 8000 Hz for 256-channel EEG data. Besides, we downsampled the original HR EEG data along the channel dimension to obtain LR counterparts at different spatial resolutions, as shown in Table I. By using electrode systems with different channel ranks, we identified and removed the channels to be estimated, resulting in LR EEG that matched the preset channel levels.
@@ IV-B1 Implement Details
The proposed STAD are implemented using the PyTorch framework, and the corresponding experiments were conducted on an Nvidia RTX A800 GPU. With the Adam optimizer, training of the model was carried out using a learning rate of $2\times 10^{-4}$, 300 epochs, and a batch size of 32. In the diffusion model framework, the forward process was set to 1000 time steps, with noise levels following sine or cosine schedules, and Gaussian noise was used for the noise type. In the multi-scale 1D convolution blocks, we employed 1D convolution layers with respective kernel sizes of 3, 5, 7, and 9. In the diffusion Transformer block, the number of attention heads was set to 16, and the hidden dimension was 64.
@@ IV-B2 Metrics
To evaluate the quality of the synthetic EEG signals, we employed four quantitative metrics: Pearson Correlation Coefficient (PCC), Normalized Mean Squared Error (NMSE), Signal-to-Noise Ratio (SNR), and Mean Absolute Error (MAE). These metrics offer a thorough understanding of the accuracy and reliability of the synthetic signals in comparison to the real signals.
@@ IV-C2 Comparison With Different scaling factors
To evaluate the performance of the proposed STAD in generating synthetic SR EEG from LR EEG under different scaling factors, we employ four metrics for quantitative analysis of the model’s generative performance. Fig. 4 shows quantitative results for different scaling factors. As the scaling factor increases, PCC and SNR values decrease, indicating a weakening correlation between synthetic SR EEG and ground truth. Meanwhile, MAE and NMSE values increase, implying a growing difference. This suggests that the reconstruction performance of STAD diminishes at larger scaling factors due to limited channel information, making HR EEG reconstruction challenging. However, even at larger scaling factors, the performance of STAD remains comparable to that at smaller scaling factors. For example, with a scaling factor of 16, the PCC value is only 3.02% lower than at a scaling factor of 8, the MAE value differs by 1.68%, and the SNR value by 3.19%. This indicates that STAD can still adequately reconstruct SR EEG signals despite limited channel information. Overall, STAD effectively enhance LR EEG spatial resolution across different scaling factors, achieving better reconstruction at smaller scales while maintaining satisfactory performance at larger scales, demonstrating strong adaptability and robustness in handling various channel levels.
@@ IV-C3 Qualitative results of Synthetic EEG
To observe the differences between synthetic SR EEG and LR EEG at different channel levels (16, 32, 64, and 128) more intuitively, we analyze the scalp topomaps of EEG in different frequency bands. Initially, the Multitaper method is employed to extract the power spectral density (PSD) of EEG with different spatial resolutions across five frequency bands, including $\gamma$ (30-40 Hz or higher), $\beta$ (13-30 Hz), $\alpha$ (8-13 Hz), $\theta$ (4-8 Hz), and $\delta$ (0.5-4 Hz) [49]. Finally, we use the MNE tool to obtain scalp topographies in different frequency bands. Fig. 5 compares scalp topographies across different frequency bands between EEG at various channel levels and model-generated 256-channel SR EEG. Each column represents a frequency band, and each row represents a type of EEG. As spatial resolution increases, the active regions on the scalp topographies become smaller and clearer. For instance, the 16-channel LR EEG shows large, blurry active regions in both low ($\delta$, $\theta$, $\alpha$) and high-frequency bands ($\beta$, $\gamma$), whereas the 256-channel SR EEG accurately localizes active sources in high-frequency bands. The synthesized SR EEG retains the frequency characteristics of the original signal across all bands, effectively reconstructing their spatial distribution. This demonstrates the strong generalization of the model across different frequency bands and its ability to stably reconstruct EEG signals. Overall, the proposed method generates SR EEG with richer spatial detail, more accurately describing and locating active regions on the scalp, whereas LR EEG results in coarse active regions, making it challenging to pinpoint active brain areas.

### v5

@@ I Introduction
Electroencephalogram (EEG) is a economical, non-invasive neuroimaging technology with high temporal resolution, playing a vital role in cognitive neuroscience research and clinical diagnosis. Unlike other neuroimaging techniques such as positron emission tomography (PET) and functional magnetic resonance imaging (fMRI), EEG offers superior temporal resolution and can continuously monitor brain activity in natural environments. This capability substantially supports the study of dynamic cognitive processes [1]. Recently, the development of portable and wireless EEG devices has further improved the flexibility and usability of EEG, enabling its widespread application in emerging neural engineering scenarios such as epileptic seizure detection and brain-computer interface (BCI) control [2, 3, 4]. Besides, these consumer devices can realize real-time remote monitoring of patients through the Internet of Medical Things (IoMT), enhance the early diagnosis and treatment, and improve the performance of real-time applications in intelligent consumer healthcare systems [5, 6]. However, the devices typically have fewer than 64 channels, resulting in lower spatial resolution of the acquired EEG data, which severely limits its potential in clinical diagnosis and brain function imaging research [7]. Currently, high-density (HD) EEG devices improve the spatial resolution of EEG by employing more electrode quantities, thus assisting doctors in performing evaluations or diagnoses in different clinical scenarios. For example, recent studies have shown that non-invasive HD EEG devices can objectively, non-invasively, and accurately determine the epileptic focus area [8, 9, 10]. However, HD EEG devices greatly increase hardware and acquisition costs, their bulkiness and complex wiring prevent subjects from moving their heads freely, restricting their use to controlled environments such as laboratories or hospitals. In addition, due to poor wearing comfort, subjects cannot be monitored for extended periods [11]. Fig. 1 summarizes the respective advantages and limitations of different density EEG devices and reveals difficulties in acquisition and limited application scenarios for HD EEG devices.
@@
Currently, in the field of enhancing spatial resolution of EEG, the methodologies predominantly fall into two main categories: channel interpolation based on mathematical models and EEG SR reconstruction utilizing deep learning methods. Channel interpolation methods principally rely on predefined scalp surface models to reconstruct missing channel data using signals from other channels, thereby improving spatial resolution. These methods involve calculating the second-order derivatives of the interpolation function in space to explore the distribution across the electrode channels [14]. Subsequent studies have attempted more precise measurements of the spatial distances between EEG electrodes and more accurate predictions for missing channel data, for example, using inverse distance weighting based on Euclidean distances and multi-quadratic interpolation methods. However, these methods only utilize the spatial correlations among adjacent electrodes and typically struggle to effectively capture and reconstruct the high-frequency and transient components of brain electrical activity resulting in suboptimal reconstruction performance for non-stationary EEG signals.
@@
Lately, deep learning has become the mainstream method in multiple fields, including the field of Internet healthcare [15, 16, 17] and medical image processing [18, 19, 20]. Specifically, deep learning-based EEG SR technology involves training deep learning models to learn the implicit mappings between LR EEG and HR EEG, and using these trained models to reconstruct SR EEG. Notably, research has introduced EEG SR techniques based on convolution neural networks (CNN) [21], followed by studies proposing deep EEG SR frameworks using graph convolution networks (GCNs) [22] to correlate the structural and functional connections of participants’ brains [23]. Additionally, generative models offer a new learning paradigm for medical image reconstruction [24]. Existing researches have utilized approaches such as Autoencoders (AE) and Generative Adversarial Networks (GANs) [25] to enhance the spatial resolution of EEG. Although current EEG SR methods based on deep learning have made considerable advancements in enhancing EEG spatial resolution, they are still limited in effectively capturing both the temporal dynamics and spatial features of EEG, resulting in limited generalizability and practicality of the models. In addition, the enhanced spatial resolution achieved by these methods is limited, with a maximum capacity to reconstruct EEG data up to 64 channel level, which may not meet the requirements of specific clinical scenarios.
@@
In clinical diagnostic and therapeutic settings, achieving high-fidelity reconstruction of HR (256 channels) ground truth EEG presents a substantial challenge. The key is accurately depicting the dependencies between LR EEG and HR EEG and addressing the significant channel-level disparity between them. Recently, a novel generative model—diffusion models [26] have exhibited remarkable generative capabilities in computer vision, leading to a significant increase in the Artificial Intelligence Generated Content (AIGC) [27, 28]. Moreover, diffusion models have demonstrated superior image generation quality compared to GANs, while also providing advantageous features such as comprehensive distribution coverage, extensibility, and stable training targets [29]. Besides, Latent Diffusion Models (LDMs) [30] extend diffusion models into latent spaces, utilizing cross-attention mechanisms to achieve refined control in conditional generation tasks. /hlFurthermore, SynDiff [31] and DiffMDD [32] have illustrated that diffusion models are applicable to generate medical images, thus advancing the precision of disease diagnosis. Despite this, applying diffusion models to enhance EEG spatial resolution has not been thoroughly explored. Therefore, this paper aims to investigate the potential of diffusion models in achieving EEG SR, particularly in enhancing spatial resolution. Through this novel generative model, we hope to provide a novel and effective method for processing and analyzing EEG data, supporting more accurate EEG studies and applications in various neuroscience contexts.
@@ III-B Basic ideas
As Fig. 2 shows, the entire model framework consists of EEG pre-trained autoencoder, STC and MTD. During the training phase of STAD, the HR EEG ground truth is first input into the EEG pre-trained encoder to obtain the corresponding latent vectors $z_{0}$. The forward diffusion process progressively derives the state $z_{t}$ of each time step $t$. In the reverse denoising process, the STC first encodes the time series and channel spatial distribution matrix from LR EEG into the spatio-temporal representation conforming to the subject, and concatenates it with time embedding as the conditioning vector. MTD takes the conditioning vectors and noise sampled from the standard normal distribution as input, implements the conditional guidance mechanism through the cross-attention layer, and predicts the noise at each step of the reverse process. The Mean Squared Error (MSE) loss is minimized between the predicted and actual noise at each step, which guides the STC module and MTD module in performing iterative parameter updates during the training phase. During the sampling phase of STAD, the deterministic MTD precisely generates SR EEG matching the input LR EEG by progressively sampling using the spatio-temporal conditional information of LR EEG and noise sampled from the Gaussian distribution.
@@ III-C2 Spatio-temporal Condition Module
As the yellow part of Fig. 2 shows, the STC is mainly composed of a spatial position embedding layer, a 1D convolution block, and a Transformer block. More specifically, the spatial position embedding layer the electrode spatial coordinates of the input LR EEG and effectively transmits the spatial structure information to the model. Then, we partition the input time series into different patches and process them with the 1D convolution block for the initial extraction of temporal features [45]. Finally, the Transformer block concatenates the temporal features with position encoding and channel spatial embeddings to output the final spatio-temporal condition $c$. The 1D convolution block consists of a 1D convolution layer, batch normalization layer and a ReLU activation function. In the Transformer block, a multi-head self-attention (MSA) layer is followed by feed-forward networks, both of which are encapsulated with residual connections and layer normalization.
@@ III-C3 Multi-scale Transformer Denoising Module
EEG, limited by the number and placement of electrodes, struggles with spatial resolution. High-density EEG devices have shown promise in tasks like source localization and BCI. The multi-scale Transformer denoising module is designed to enhance the spatial resolution of existing EEG devices. Traditional diffusion models, which typically use UNet, are effective in image generation but limited in handling the long-sequence dependencies of time-series data like EEG. Inspired by [41], the MTD employs a Transformer backbone to more effectively extract spatio-temporal features. Additionally, multi-scale 1D convolution blocks are introduced to capture multi-scale temporal features in the reverse denoising process. This structure allows the model to process signals across different frequency ranges simultaneously [46].
@@ III-D Loss Function
where $\epsilon_{\theta}$ is the proposed multi-scale Transformer denoising model and $\tau_{\theta}$ represents the proposed STC module. Algorithm 1 presents the entire training procedure along with the final loss function.
@@
      Perform: Extract spatio-temporal features of LR EEG $x$ via STC $\tau_{\theta}$ ;
      $c=\tau_{\theta}\left(x\right)$ ;
@@ IV-A Dataset and Preprocessing
We employed the MNE software for preprocessing of the raw data. In this work, we applied high-pass filtering to the continuous EEG time series to remove low-frequency irrelevant information. To eliminate the large line noise caused by direct current, we used notch filtering at 50, 100, 150, and 200 Hz. Subsequently, we segmented the raw data into epochs of fixed duration (350 ms) based on the provided annotation events in the dataset. Moreover, the dataset was recorded using the EGI NA-400 amplifier (Electrical Geodesics) [51] at a sampling rate of 8000 Hz for 256-channel EEG data. Besides, we downsampled the original HR EEG data along the channel dimension to obtain LR counterparts at different spatial resolutions. As shown in Table I, we set the SR scaling factors to 2, 4, 8, and 16. Specifically, each scaling factor corresponds to a specific number of channels from the EGI standard electrode system, thereby simulating the acquisition of EEG data with varying densities in real-world scenarios. Following the preprocessing steps, the paried dataset comprises a total of 23,930 samples. , with 80%, or 19,144 samples, allocated to the training set for model learning, 10%, or 2,393 samples, designated as the testing set to assess the model’s generalization capabilities, and the remaining 10%, also 2,393 samples, serving as the validation set for evaluation of model performance.
@@ IV-B1 Implement Details
The proposed STAD are implemented using the PyTorch framework, and the corresponding experiments were conducted on an Nvidia RTX A800 GPU. With the Adam optimizer, training of the model was carried out using a learning rate of $2\times 10^{-4}$, 300 epochs, and a batch size of 32. In the diffusion model framework, the time step of the forward process was set to 1000, during which the noise levels were modulated by sine or cosine schedules, and Gaussian noise was used for the noise type. In the multi-scale 1D convolution blocks, we employed 1D convolution layers with respective kernel sizes of 3, 5, 7, and 9. In the diffusion Transformer block, the number of attention heads was set to 16, and the hidden dimension was 64.
@@ IV-B2 Metrics
To evaluate the quality of synthetic SR EEG, we employed four quantitative metrics: Pearson Correlation Coefficient (PCC), Normalized Mean Squared Error (NMSE), Signal-to-Noise Ratio (SNR), and Mean Absolute Error (MAE). These metrics offer a thorough understanding of the similarity and reliability of the synthetic signals in comparison to the real signals.
@@ IV-C2 Comparison with Different Scaling Factors
To evaluate the performance of the proposed STAD in generating synthetic SR EEG from LR EEG under different scaling factors, we employ four metrics for quantitative analysis of the model’s generative performance. Fig. 4 shows quantitative results for different scaling factors. As the scaling factor increases, PCC and SNR values decrease, indicating a weakening correlation between synthetic SR EEG and ground truth. Meanwhile, MAE and NMSE values increase, implying a growing difference. The performance degradation of the model at high scaling factors can be attributed to two main reasons: Firstly, when larger scaling factors are used, the difference in the number of channels between LR EEG and HR EEG becomes more pronounced. Secondly, with fewer channels in the LR EEG input, the signal lacks sufficient temporal and spatial details, which makes it challenging for the model to recover the correct spatial structure and dynamic patterns during the reconstruction process.
However, even at larger scaling factors, the performance of STAD remains comparable to that at smaller scaling factors. For example, with a scaling factor of 16, the PCC value is only 3.02% lower than at a scaling factor of 8, the MAE value differs by 1.68%, and the SNR value by 3.19%. This indicates that STAD can still adequately reconstruct SR EEG signals despite limited channel information. Overall, STAD effectively enhance LR EEG spatial resolution across different scaling factors, achieving better reconstruction at smaller scales while maintaining satisfactory performance at larger scales, demonstrating strong adaptability and robustness in handling various channel levels.
@@ IV-C3 Qualitative Results of Synthetic EEG
To observe the differences between synthetic SR EEG and LR EEG at different channel levels (16, 32, 64, and 128) more intuitively, we analyze the scalp topomaps of EEG in different frequency bands. Initially, the Multitaper method is employed to extract the power spectral density (PSD) of EEG with different spatial resolutions across five frequency bands, including $\gamma$ (30-40 Hz or higher), $\beta$ (13-30 Hz), $\alpha$ (8-13 Hz), $\theta$ (4-8 Hz), and $\delta$ (0.5-4 Hz) [52]. Finally, we use the MNE tool to obtain scalp topographies in different frequency bands. Fig. 5 compares scalp topographies across different frequency bands between EEG at various channel levels and synthetic 256-channel SR EEG. Each column represents a frequency band, and each row represents a type of EEG. As spatial resolution increases, the active regions on the scalp topographies become smaller and clearer. For instance, the 16-channel LR EEG shows large, blurry active regions in both low ($\delta$, $\theta$, $\alpha$) and high-frequency bands ($\beta$, $\gamma$), whereas the 256-channel SR EEG accurately localizes active sources in high-frequency bands. The synthesized SR EEG retains the frequency characteristics of the original signal across all bands, effectively reconstructing their spatial distribution. Furthermore, we observed that the synthetic SR EEG is able to retain the frequency band characteristics of the original EEG signal. This demonstrates the strong generalization of the model across different frequency bands and its ability to stably reconstruct EEG signals. In general, the proposed method generates SR EEG with richer spatial detail, more accurately describing and locating active regions on the scalp, while LR EEG results in coarse active regions, making it challenging to pinpoint active areas of the brain.